zarkam21 (zarkam21):

What is the vertex of the parabola given by the equation below? x=-4(y-3)^2+2

zarkam21 (zarkam21):

Is it -4?

OpenStudy (mathmate):

Hint: The conventional equation of a parabola in vertex form is: \(y=a(x-h)^2+k\), where the vertex is at (h,k). If a is positive, parabola opens upwards. However, we can turn the parabola 90 degrees, i.e. let the parabola sit sideways, opens left/right, by exchanging x and y to give \(x=a(y-k)^2+h\), which makes the vertex at again (h,k). if a is positive, parabola opens to the right. by comparing the given equation with the standard forms, the solution can be obtained. Note: a vertex is defined by an ordered pair, like (4,-2) and not a single number.

zarkam21 (zarkam21):

(3,-2)?

OpenStudy (phi):

Did you follow mathmate's post? the vertex is an (x,y) point. it is (h,k) where h and k match up with x= a (y − k)^2+h x=-4 (y - 3)^2+2 what number matches up with "h" ? and what number matches up with k ?

zarkam21 (zarkam21):

With h 2 and with k -3

OpenStudy (phi):

yes, h is 2 but